# Multi-Objective-Guided Discrete Flow Matching for Controllable Biological Sequence Design

### v1

@@ Abstract
Designing biological sequences that satisfy multiple, often conflicting, functional and biophysical criteria remains a central challenge in biomolecule engineering. While discrete flow matching models have recently shown promise for efficient sampling in high-dimensional sequence spaces, existing approaches address only single objectives or require continuous embeddings that can distort discrete distributions. We present Multi-Objective-Guided Discrete Flow Matching (MOG-DFM), a general framework to steer any pretrained discrete-time flow matching generator toward Pareto-efficient trade-offs across multiple scalar objectives. At each sampling step, MOG-DFM computes a hybrid rank-directional score for candidate transitions and applies an adaptive hypercone filter to enforce consistent multi-objective progression. We also trained two unconditional discrete flow matching models, PepDFM for diverse peptide generation and EnhancerDFM for functional enhancer DNA generation, as base generation models for MOG-DFM. We demonstrate MOG-DFM’s effectiveness in generating peptide binders optimized across five properties (hemolysis, non-fouling, solubility, half-life, and binding affinity), and in designing DNA sequences with specific enhancer classes and DNA shapes. In total, MOG-DFM proves to be a powerful tool for multi-property-guided biomolecule sequence design.
@@ 2 Discrete Flow Matching
In the discrete setting, we consider data $x=(x_{1},\dots,x_{d})$ taking values in a finite state space $S=\mathcal{T}^{d}$, where $\mathcal{T}=[K]=\{1,2,\dots,K\}$ is called the vocabulary. We model a continuous-time Markov chain (CTMC) $\{X_{t}\}_{t\in[0,1]}$ whose time-dependent transition rates $u_{t}(y,x)$ transport probability mass from an initial distribution $p_{0}$ to a target distribution $p_{1}$ [25]. The marginal probability at time $t$ is denoted $p_{t}(x)$, and its evolution is governed by the Kolmogorov forward equation
@@
In practice, we can further parametrize the velocity field using a mixture path. Specifically, one defines a mixture path with scheduler $\kappa_{t}\in[0,1]$ so that each coordinate $X_{t}^{(i)}$ equals $x_{0}^{(i)}$ or $x_{1}^{(i)}$ with probabilities $1-\kappa_{t}$ and $\kappa_{t}$ respectively. The mixture marginal velocity is then obtained by averaging the conditional rates over the posterior of $(x_{0},x_{1})$ given $X_{t}=x$, yielding
@@
where $\dot{\kappa}_{t}$ denotes the time derivative of ${\kappa}_{t}$. Therefore, the aim of discrete flow matching model training, which is to learn the velocity field $u_{t}^{i}(y^{i},x)$, now equals to learning the marginal posterior $p^{i}_{1\mid t}(x_{1}^{i}\mid x)$. In this case, we can set the Bregman divergence to the generalized KL comparing general vectors $u,v\in\mathbb{R}_{\geq 0}^{m}$,
@@ 3 Multi-Objective Guided Discrete Flow Matching
MOG-DFM (Multi-Objective Guided Discrete Flow Matching) operates under the same setting as discrete flow matching described in the previous section. Suppose we have a pre-trained discrete flow matching model that defines a CTMC with factorized velocity field $u_{t}^{i}(y^{i},x)$ , which transports probability mass from an initial distribution $p_{0}$ to the unknown target distribution via mixture path parametrization. In addition, we assume access to $N$ pre-trained scalar score functions $s_{n}:\mathcal{S}\to\mathbb{R},\text{where}\ n=1,\dots,N$, that assign objective scores to any sequence. Our aim is to generate novel sequences $x_{1}\in\mathcal{S}$ whose objective vectors $\bigl{(}s_{1}(x_{1}),\,s_{2}(x_{1}),\,\dots,s_{N}(x_{1})\bigr{)}$ lie near the Pareto front (not guaranteed to be Pareto optimal)
@@ 3.1 Step 0: Initialization and Weight Vector Generation
MOG-DFM begins by initializing the generative process at time $t=0$ by sampling an initial sequence $x_{0}$ uniformly from the discrete state space $\mathcal{S}=[K]^{d}$. To steer the generation towards diverse Pareto-efficient solutions, we introduce a set of weight vectors $\{\omega^{k}\}_{k=1}^{M}$ that uniformly cover the $N$-dimensional Pareto Front. Intuitively, each $\omega$ encodes a particular trade-off among the $N$ objectives, so sampling different $\omega$ promotes exploration of distinct regions of the Pareto front. Concretely, we construct these vectors via the Das–Dennis simplex lattice with $H$ subdivisions, yielding components
@@
and then draw one $\omega$ randomly before the following steps. This defines one direction we want to optimize in the Pareto Flow for the current run. The following three steps will then be performed in each iteration. We set the number of total iterations to be $T$.
@@ 3.2 Step 1: Guided Transition Scoring
We first randomly select one position $i$ on the sequence so that we will update the token on this position during the current iteration. At each intermediate state $x_{t}$ and selected position $i$, each possible candidate transition $y^{i}\neq x^{i}$ is scored by combining local improvement measures with global directional alignment. The normalized rank score captures how much each individual objective improves relative to other possible token replacements, thereby encouraging exploration of promising local moves; formally, for each objective $n$ we compute
@@
measures the alignment of the multi-objective improvement vector $\Delta\mathbf{s}$ with the chosen weight vector $\omega$, ensuring that transitions not only improve individual objectives but collectively move toward the desired trade-off direction. By z-score normalizing both components and combining them as
@@
we balance rank-based exploration against direction-guided exploitation with $\lambda>0$. An importance vector $\mathbf{I}=[i_{1},\dots,i_{N}]$ is used to normalize the improvement values for each objective. Finally, we re-weight the original factorized velocity field from the pre-trained discrete flow matching model:
$u_{\mathrm{guided},t}^{\,i}(y^{i},x\mid\omega)=\begin{cases}\beta\,u_{t}^{\,i}(y^{i},x)\,\exp\bigl{(}\Delta S(y^{i},x,\omega)\bigr{)},\quad y^{i}\neq x^{i}\\ -\sum_{y^{i}\neq x^{i}}u_{\mathrm{guided},t}^{\,i}(y^{i},x\mid\omega),\quad y^{i}=x^{i}\end{cases}$ (15)
@@ 3.3 Step 2: Adaptive Hypercone Filtering
There are two degenerate cases that may lead to empty $Y^{i}$: If every $\alpha^{i}\geq\pi$, indicating that all possible transitions decrease performance, we will perform a self-transition and retain the current state; if there exist some $\alpha^{i}<\pi$ but none lie within the cone (i.e., $\Phi$ is temporarily too small), we still advance by choosing the best-aligned candidate
@@ 3.4 Step 3: Euler Sampling
Here, $h=1/T$ is the step size in the time interval, $X_{t}$ and $X_{t+h}$ denotes the current state and the next state respectively. In practice, one draws a uniform random number $r\in[0,1]$: if $r\leq 1-\exp(-h\,R_{t}^{\,i}(x))$, $x^{i}$ will transition to the best selected candidate; otherwise we retain $x^{i}$.
After performing from step 1 to step 3 for $T$ iterations, we end with the final sample $x_{1}$ whose score vectors have been steered close to the Pareto Front, with all objectives optimized.

### v2

@@ Abstract
Designing biological sequences that satisfy multiple, often conflicting, functional and biophysical criteria remains a central challenge in biomolecule engineering. While discrete flow matching models have recently shown promise for efficient sampling in high-dimensional sequence spaces, existing approaches address only single objectives or require continuous embeddings that can distort discrete distributions. We present Multi-Objective-Guided Discrete Flow Matching (MOG-DFM), a general framework to steer any pretrained discrete flow matching generator toward Pareto-efficient trade-offs across multiple scalar objectives. At each sampling step, MOG-DFM computes a hybrid rank-directional score for candidate transitions and applies an adaptive hypercone filter to enforce consistent multi-objective progression. We also trained two unconditional discrete flow matching models, PepDFM for diverse peptide generation and EnhancerDFM for functional enhancer DNA generation, as base generation models for MOG-DFM. We demonstrate MOG-DFM’s effectiveness in generating peptide binders optimized across five properties (hemolysis, non-fouling, solubility, half-life, and binding affinity), and in designing DNA sequences with specific enhancer classes and DNA shapes. In total, MOG-DFM proves to be a powerful tool for multi-property-guided biomolecule sequence design.
@@ 2 Discrete Flow Matching
In the discrete setting, we consider data $x=(x_{1},\dots,x_{d})$ taking values in a finite state space $S=\mathcal{T}^{d}$, where $\mathcal{T}=[K]=\{1,2,\dots,K\}$ is called the vocabulary. We model a continuous-time Markov chain (CTMC) $\{X_{t}\}_{t\in[0,1]}$ whose time-dependent transition rates $u_{t}(y,x)$ transport the probability mass from an initial distribution $p_{0}$ to a target distribution $p_{1}$ [25]. The marginal probability at time $t$ is denoted $p_{t}(x)$, and its evolution is governed by the Kolmogorov forward equation
@@
In practice, we can further parameterize the velocity field using a mixture path. Specifically, a mixture path is defined with scheduler $\kappa_{t}\in[0,1]$ so that each coordinate $X_{t}^{i}$ equals $x_{0}^{i}$ or $x_{1}^{i}$ with probabilities $1-\kappa_{t}$ and $\kappa_{t}$, respectively. The mixture marginal velocity is then obtained by averaging the conditional rates over the posterior of $(x_{0},x_{1})$ given $X_{t}=x$, yielding
@@
where $\dot{\kappa}_{t}$ denotes the time derivative of ${\kappa}_{t}$. Therefore, the aim of discrete flow matching model training, which is to learn the velocity field $u_{t}^{i}(y^{i},x)$, is now equivalent to learning the marginal posterior $p^{i}_{1\mid t}(x_{1}^{i}\mid x)$. In this case, we can set the Bregman divergence to the generalized KL comparing general vectors $u,v\in\mathbb{R}_{\geq 0}^{m}$,
@@ 3 Multi-Objective Guided Discrete Flow Matching
MOG-DFM (Multi-Objective Guided Discrete Flow Matching) operates under the same setting as discrete flow matching described in the previous section. Suppose we have a pre-trained discrete flow matching model that defines a CTMC with a factorized velocity field $u_{t}^{i}(y^{i},x)$ , which transports probability mass from an initial distribution $p_{0}$ to the unknown target distribution via mixture path parametrization. In addition, we assume access to $N$ pre-trained scalar score functions $s_{n}:\mathcal{S}\to\mathbb{R},\text{where}\ n=1,\dots,N$, that assign objective scores to any sequence. Our aim is to generate novel sequences $x_{1}\in\mathcal{S}$ whose objective vectors $\bigl{(}s_{1}(x_{1}),\,s_{2}(x_{1}),\,\dots,s_{N}(x_{1})\bigr{)}$ lie near the Pareto front (not guaranteed to be Pareto optimal)
@@ 3.1 Step 0: Initialization and Weight Vector Generation
MOG-DFM begins by initializing the generative process at time $t=0$ by sampling an initial sequence $x_{0}$ uniformly from the discrete state space $\mathcal{S}=[K]^{d}$. To steer the generation towards diverse Pareto-efficient solutions, we introduce a set of $M$ weight vectors $\{\omega^{k}\}_{k=1}^{M}$, where $\omega\in\mathbb{R}^{N}$, that uniformly cover the $N$-dimensional Pareto front. Intuitively, each $\omega$ encodes a particular trade-off among the $N$ objectives, so sampling different $\omega$ promotes exploration of distinct regions of the Pareto front. We construct these vectors via the Das–Dennis simplex lattice with $H$ subdivisions, yielding components
@@
and then draw one $\omega$ randomly before the following steps. This defines one direction we want to optimize toward the Pareto front in the current run. The following three steps will then be performed in each iteration. We set the number of total iterations to $T$.
@@ 3.2 Step 1: Guided Transition Scoring
We first randomly select one position $i$ in the sequence that will be updated during the current iteration. At each intermediate state $x_{t}$ and selected position $i$, each possible candidate transition $y^{i}\neq x^{i}$ is scored by combining local improvement measures with global directional alignment. The normalized rank score captures how much each individual objective improves relative to other possible token replacements, thereby encouraging exploration of promising local moves. Formally, for each objective $n$ we compute
@@
measures the alignment of the multi-objective improvement vector $\Delta\mathbf{s}$ with the chosen weight vector $\omega$, ensuring that transitions not only improve individual objectives but collectively move toward the desired trade-off direction. To balance rank-based exploration against direction-guided exploitation, we z-score normalize both components and combine them as follows
@@
where $\lambda>0$ is a tunable hyperparameter. An importance vector $\mathbf{I}=[i_{1},\dots,i_{N}]$ is used to normalize the improvement values for each objective. Finally, we re-weight the original factorized velocity field from the pre-trained discrete flow matching model:
$u_{\mathrm{guided},t}^{\,i}(y^{i},x\mid\omega)=\begin{cases}\beta\,u_{t}^{\,i}(y^{i},x)\,\exp\bigl{(}\Delta S(y^{i},x,\omega)\bigr{)}&\quad y^{i}\neq x^{i}\\ -\sum_{y^{i}\neq x^{i}}u_{\mathrm{guided},t}^{\,i}(y^{i},x\mid\omega)&\quad y^{i}=x^{i}\end{cases}$ (15)
@@ 3.3 Step 2: Adaptive Hypercone Filtering
There are two degenerate cases that can lead to empty $Y^{i}$: (1) if every $\alpha^{i}\geq\pi$, indicating that all possible transitions decrease performance, we will perform a self-transition and retain the current state; (2) if there exist some $\alpha^{i}<\pi$ but none lie within the cone (i.e., $\Phi$ is temporarily too small), we still advance by choosing the best-aligned candidate
@@ 3.4 Step 3: Euler Sampling
Here, $h=1/T$ is the step size in the time interval, $X_{t}$ and $X_{t+h}$ denotes the current state and the next state respectively. In practice, one draws a uniform random number $r\in[0,1]$: if $r\leq 1-\exp(-h\,R_{t}^{\,i}(x))$, $x^{i}$ will transition to the best selected candidate; otherwise, we retain $x^{i}$.
After performing step 1 to step 3 for $T$ iterations, we end with the final sample $x_{1}$ whose score vectors have been steered close to the Pareto front, with all objectives optimized.